PROGRESS TOWARDS THE ROTATIONAL SPECTRUM OF H_{5}^{+} AND ITS ISOTOPOLOGUES

\underline{BRETT A. MCGUIRE}, YIMIN WANG, JOEL M. BOWMAN, AND SUSANNA L. WIDICUS WEAVER, Department of Chemistry, Emory University, Atlanta, GA 30033.

The reaction of H_{3}^{+} with H_{2}, arguably the most common bimolecular reaction in the universe, proceeds through the H_{5}^{+} collisional complex. This reaction, and consequently H_{5}^{+}, greatly influence the chemical and physical processes in the interstellar medium, playing crucial roles in such varied processes as isotopic fractionation and the formation of complex organic molecules. A thorough understanding of the role of H_{5}^{+} in interstellar chemistry is contingent upon its definitive astronomical detection, necessitating the acquisition of a laboratory rotational spectrum. Rotationally-resolved spectra of H_{5}^{+} in the terahertz region have not yet been observed experimentally. The prediction of this spectrum based on a highlevel theoretical study is therefore an important first step to guide experiment. The highly fluxional nature of H_{5}^{+} presents major challenges for theory, especially for the pure rotational spectrum due to the difficulties in determining an accurate dipole moment from a correct description of the highly delocalized zero-point wavefunction. We have now completed this work using the most recent potential energy and dipole moment surfaces for H_{5}^{+} and its isotopologues DH_{4}^{+}, D_{2}H_{3}^{+}, D_{3}H_{2}^{+}, D_{4}H^{+}, and D_{5}^{+}. Pure rotational spectra have been predicted for these species based on the optimized minimum-energy geometries and the zero-point averaged dipole moments calculated from our potential energy surface. We will discuss the implications of these results for the detection of each ion’s rotational spectrum, show preliminary predictions of the rotational spectrum for those species possessing permanent dipole moments, and comment on the degree of expected spectral splitting arising from internal motion. Finally, we will report on progress toward the laboratory spectroscopic investigation of these species in the terahertz region.